# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 3.1 The NPDo Ansatz
    When $f(P)$ is right-unitarily invariant, it suffices to take $Q=I_{k}$ and $\widetilde{P}=\widehat{P}$ because $f(\widehat{P})=f(\widetilde{P})$ regardless of $Q$. Introducing subset $\mathbb{P}$ of $\mathbb{O}^{n\times k}$ and judiciously choosing $Q$ are for generality in order to deal with the case when $f(P)$ is not right-unitarily invariant, e.g., the one in Example 3.1 and those from Table 1 in section 1 that involve $D$ or $D_{i}$. Throughout this paper, we will assume that $\mathbb{P}$ is as inclusive as necessary to allow our proving arguments to go through. In particular, at the minimum, $\mathbb{P}$ should contain one or more maximizers of the associated optimization problem (1.1).
@@
    For computational purposes, it is necessary to have an efficient way to construct $Q$ in the ansatz. That is often the case when it comes to common concrete objective functions $f$ that are in use today. In our later development, either a proper $\mathbb{P}$ can maximally increase the value of objective function $f$, e.g., when $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ for $m\geq 1$ is involved, or we have to have it for our theoretical proofs to go through. In fact for $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$, we may take $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$, and let $Q\in\mathbb{O}^{k\times k}$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$ to ensure $\widetilde{P}^{\operatorname{T}}D=Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$. As a consequence, $(\widetilde{P}^{\operatorname{T}}D)^{m}\succeq 0$ and $\|(\widetilde{P}^{\operatorname{T}}D)^{m}\|_{\operatorname{tr}}=\operatorname{tr}((\widetilde{P}^{\operatorname{T}}D)^{m})$ by Lemma B.9 and hence an orthogonal polar factor $Q$ of $\widehat{P}^{\operatorname{T}}D$ maximizes $\operatorname{tr}([(\widehat{P}Z)^{\operatorname{T}}D]^{m})$ over $Z\in\mathbb{O}^{k\times k}$. Calculating this $Q$ via the SVD of $\widehat{P}^{\operatorname{T}}D\in\mathbb{R}^{k\times k}$ is efficient since $k$ is usually small (usually no more than a couple of hundreds usually).
@@
    It is tempting to stipulate $f(\widehat{P})\geq f(P)+\omega\eta$, but that is either false or just hard to prove, e.g., for the one in Example 3.1. Often in our algorithms to solve (1.1) iteratively, with $P$ being the current approximate maximizer, assuming the NPDo Ansatz, we naturally compute $\widehat{P}$ that maximizes $\operatorname{tr}(X^{\operatorname{T}}\mathscr{H}(P))$ over $X\in\mathbb{O}^{n\times k}$. With that $\widehat{P}$, settling whether $f(\widehat{P})\geq f(P)+\omega\eta$ or not can be a hard or even impossible task, for example, in Example 3.1 it is not clear if $f(\widehat{P})\geq f(P)+\omega\eta$ at all.
@@
As to the validity of the NPDo Ansatz on the objective functions in Table 1, it holds for all, except for those that involve quotients, under reasonable conditions on the constant matrices and function $\phi$. Table 2 details conditions under which the NPDo Ansatz holds, where the last column points to the places for justifications. We point out that we can take $\mathbb{P}=\mathbb{O}^{n\times k}$, $Q=I_{k}$, and $\omega=1$ for all in Table 2 but judicious choices of $\mathbb{P}$ and $Q$ can increase the values of objective functions more than $\omega\eta$ as stipulated by the NPDo Ansatz for MBSub and SumCT [66, Theorem 5.2].
@@
Suppose that the NPDo Ansatz holds, and let $P_{*}\in\mathbb{P}$ be a maximizer of (1.1). Then (2.3) holds for $P=P_{*}$ and $\Lambda=\Lambda_{*}:=P_{*}^{\operatorname{T}}\mathscr{H}(P_{*})\succeq 0$.
@@ 3.2 SCF Iteration and Convergence
3:     $\widehat{P}^{(i)}=U_{i}V_{i}^{\operatorname{T}}\in\mathbb{O}^{n\times k}$, an orthogonal polar factor of $\mathscr{H}(P^{(i)})$;
4:     calculate $Q_{i}\in\mathbb{O}^{k\times k}$ whose existence is stipulated by the NPDo Ansatz and let $P^{(i+1)}=\widehat{P}^{(i)}Q_{i}\in\mathbb{P}$;
@@
Next, we will state our convergence theorems for Algorithm 1 under the assumption of the NPDo Ansatz. It is shown that as the SCF iteration progresses, the value of the objective function monotonically increases, any accumulation point of the generated approximation sequence satisfies the necessary conditions in Theorem 3.1 for a global maximizer, and under certain conditions, the accumulation point can be proved to be the limit point of the entire approximation sequence. In short, the NPDo approach is guaranteed to work.
@@
Both Theorem 3.2(b,c) have useful consequences. As a corollary of Theorem 3.2(b), we find that the NPDo Ansatz is a sufficient condition for NPDo (2.3) to have a solution because there always exists an accumulation point $P_{*}$ of the sequence $\{P^{(i)}\}_{i=0}^{\infty}$ in $\mathbb{O}^{n\times k}$.
@@
Under the NPDo Ansatz, NPDo (2.3) is solvable, i.e., there exists $P\in\mathbb{O}^{n\times k}$ such that $\Lambda=P^{\operatorname{T}}\mathscr{H}(P)\succeq 0$ and (2.3) holds.
As a corollary of Theorem 3.2(c), if $\sigma_{\min}(\mathscr{H}(P^{(i)})$ is eventually bounded below away from 00 uniformly [^2] By which we mean that there exist a constant $\tau>0$ and an integer $K$ such that $\sigma_{\min}(\mathscr{H}(P^{(i)})\geq\tau$ for all $i\geq K$., then
@@
    $\operatorname{rank}(\mathscr{H}(P_{*}Q))=k\quad\mbox{for any $Q\in\mathbb{O}^{k\times k}$},$ (3.7)
@@
    and if $f(P_{*})>f(P)$ for any $P\neq P_{*}$ and ${\cal R}(P)={\cal R}(P_{*})$, i.e., $f(P)$ has unique maximizer in the orbit $\{P_{*}Q\,:\,Q\in\mathbb{O}^{k\times k}\}$, then the entire sequence $\{P^{(i)}\}_{i=0}^{\infty}$ converges to $P_{*}$.
@@
In the case when objective function $f$ is right-unitarily invariant, Theorem 3.3(c) is clearly not applicable. In such a case, computing ${\cal R}(P_{*})$ is the ultimate goal because each maximizer $P_{*}$ is really a representative from the orbit $\{P_{*}Q\,:\,Q\in\mathbb{O}^{k\times k}\}$. Given $Q\in\mathbb{R}^{n\times k}$, let $g(P)=f(PQ)$. It can be verified that
@@
Thus if $f$ is right-unitarily invariant, then for $Q\in\mathbb{O}^{n\times k}$ we have $\mathscr{H}(PQ)=\mathscr{H}(P)\,Q$ and as a result, condition (3.7) is equivalently to $\operatorname{rank}(\mathscr{H}(P_{*}))=k$.
@@ 3.3 Acceleration by LOCG and Convergence
Without loss of generality, let $P^{(-1)}\in\mathbb{O}^{n\times k}$ be the approximate maximizer of (1.1) from the very previous iterative step, and $P\in\mathbb{O}^{n\times k}$ the current approximate maximizer. We are now looking for the next approximate maximizer $P^{(1)}$, along the line of LOCG, according to
$P^{(1)}=\arg\max_{Y\in\mathbb{O}^{n\times k}}f(Y),\,\,\mbox{s.t.}\,\,{\cal R}(Y)\subseteq{\cal R}([P,\mathscr{R}(P),P^{(-1)}]),$ (3.9)
@@
$\mathscr{R}(P):=\operatorname{grad}f_{|{{\mathbb{O}}^{n\times k}}}(P)=\mathscr{H}(P)-P\cdot\frac{1}{2}\Big{[}P^{\operatorname{T}}\mathscr{H}(P)+\mathscr{H}(P)^{\operatorname{T}}P\Big{]}$ (3.10)
@@
We still have to numerically solve (3.9). For that purpose, let $W\in\mathbb{O}^{n\times m}$ be an orthonormal basis matrix of subspace ${\cal R}([P,\mathscr{R}(P),P^{(-1)}])$. Generically, $m=3k$ but $m<3k$ can happen. It can be implemented by the Gram-Schmidt orthogonalization process, starting with orthogonalizing the columns of $\mathscr{R}(P)$ against $P$ since $P\in\mathbb{O}^{n\times k}$ already. In MATLAB, to fully take advantage of its optimized functions, we simply set $W=[\mathscr{R}(P),P^{(-1)}]$ (or $W=\mathscr{R}(P)$ for the first iteration) and then do
@@
$Y=WZ\quad\mbox{for $Z\in\mathbb{O}^{m\times k}$}.$ (3.11a)
@@
$Z_{\operatorname*{opt}}=\arg\max_{Z\in\mathbb{O}^{m\times k}}\widetilde{f}(Z)\quad\mbox{with}\quad\widetilde{f}(Z):=f(WZ),$ (3.11b)
@@
~$\widetilde{\mathscr{H}}(Z)=Z\widetilde{\Lambda}\quad\mbox{with}\quad\widetilde{\Lambda}^{\operatorname{T}}=\widetilde{\Lambda}\in\mathbb{R}^{k\times k},\quad Z\in\mathbb{O}^{m\times k}.$ (3.12b)
@@
Suppose that the NPDo Ansatz holds for $f$, and let $\mathbb{Z}:=W^{\operatorname{T}}\mathbb{P}\subseteq\mathbb{O}^{m\times k}$. If $W\mathbb{Z}\subseteq\mathbb{P}$, then the NPDo Ansatz holds for $\widetilde{f}$ defined in (3.11b).
@@
Let $Z\in\mathbb{Z}$ and $\widehat{Z}\in\mathbb{O}^{m\times k}$ such that
@@
Set $P=WZ\in\mathbb{P}$ (because of $W\mathbb{Z}\subseteq\mathbb{P}$) and $\widehat{P}=W\widehat{Z}\in\mathbb{O}^{n\times k}$. Noticing that ~$\widetilde{\mathscr{H}}(Z))=W^{\operatorname{T}}\mathscr{H}(WZ)$, we have (3.1) from (3.13). By the NPDo Ansatz for $f$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q=W(\widehat{Z}Q)=:W\widetilde{Z}\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$. Hence,
@@
3:     compute $W\in\mathbb{O}^{n\times m}$ such that ${\cal R}(W)={\cal R}([P^{(i)},\mathscr{R}(P^{(i)}),P^{(i-1)}])$ such that $P^{(i)}$ occupies the first $k$ columns of $W$;
@@
We will perform a convergence analysis for Algorithm 2, considering an ideal situation that at its Line 4, $Z_{\operatorname*{opt}}$ is computed to be an exact maximizer of (3.11) for simplicity. We point out that the seemingly ideal situation is not completely unrealistic. In actual computation, as we explained in Remark 3.2(ii), the computed $Z_{\operatorname*{opt}}$ should be sufficiently more accurate as an approximation solution for (3.11) than $P^{(i)}$ as an approximation solution for the original problem (1.1) at that moment.
Before we state our main results, let us look at whether $f(P^{(i_{0}+1)})=f(P^{(i_{0})})$ can possibly happen for some $i_{0}$ in Algorithm 2. Suppose that is case. Then there will be no increase in the objective value by $P^{(i_{0}+1)}$ at Line 5 for (3.11). By (3.12), we conclude that, for $i=i_{0}$ at Line 4,
$W^{\operatorname{T}}\mathscr{H}(WZ^{(0)})=Z^{(0)}\Lambda\quad\mbox{for some $\Lambda\succeq 0$}.$
Equivalently, $0=W^{\operatorname{T}}\big{[}\mathscr{H}(WZ^{(0)})-WZ^{(0)}\Lambda\big{]}=W^{\operatorname{T}}\mathscr{R}(P^{(i_{0})})$ because $WZ^{(0)}=P^{(i_{0})}$. Since ${\cal R}(W)\supseteq{\cal R}(\mathscr{R}(P^{(i_{0})}))$, we have $\mathscr{R}(P^{(i_{0})})=0$. This implies $P^{(i_{0})}$ is a KKT point of (1.1) and satisfies the necessary conditions in Theorem 3.1 for a global maximizer. This in general is too good a scenario and almost never happens in actual computation. For that reason, in our analysis below, we will exclude such a situation from consideration.

### v4

@@ 3.1 The NPDo Ansatz
    When $f(P)$ is right-unitarily invariant, it suffices to take $Q=I_{k}$ and $\widetilde{P}=\widehat{P}$ because $f(\widehat{P})=f(\widetilde{P})$ regardless of $Q$. Introducing subset $\mathbb{P}$ of ${\rm St}(k,n)$ and judiciously choosing $Q$ are for generality in order to deal with the case when $f(P)$ is not right-unitarily invariant, e.g., the one in Example 3.1 and those from Table 1 in section 1 that involve $D$ or $D_{i}$. Throughout this paper, we will assume that $\mathbb{P}$ is as inclusive as necessary to allow our proving arguments to go through. In particular, at the minimum, $\mathbb{P}$ should contain one or more maximizers of the associated optimization problem (1.1).
@@
    For computational purposes, it is necessary to have an efficient way to construct $Q$ in the ansatz. That is often the case when it comes to common concrete objective functions $f$ that are in use today. In our later development, either a proper $\mathbb{P}$ can maximally increase the value of objective function $f$, e.g., when $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ for $m\geq 1$ is involved, or we have to have it for our theoretical proofs to go through. In fact for $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$, we may take $\mathbb{P}={\rm St}(k,n)_{D+}$, and let $Q\in{\rm St}(k,k)$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$ to ensure $\widetilde{P}^{\operatorname{T}}D=Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$. As a consequence, $(\widetilde{P}^{\operatorname{T}}D)^{m}\succeq 0$ and $\|(\widetilde{P}^{\operatorname{T}}D)^{m}\|_{\operatorname{tr}}=\operatorname{tr}((\widetilde{P}^{\operatorname{T}}D)^{m})$ by Lemma B.9 and hence an orthogonal polar factor $Q$ of $\widehat{P}^{\operatorname{T}}D$ maximizes $\operatorname{tr}([(\widehat{P}Z)^{\operatorname{T}}D]^{m})$ over $Z\in{\rm St}(k,k)$. Calculating this $Q$ via the SVD of $\widehat{P}^{\operatorname{T}}D\in\mathbb{R}^{k\times k}$ is efficient since $k$ is usually small (in the tens or no more than a couple of hundreds).
@@
    It is tempting to stipulate $f(\widehat{P})\geq f(P)+\omega\eta$, but that is either false or just hard to prove, e.g., for the one in Example 3.1. Often in our algorithms to solve (1.1) iteratively, with $P$ being the current approximate maximizer, assuming the NPDo Ansatz, we naturally compute $\widehat{P}$ that maximizes $\operatorname{tr}(X^{\operatorname{T}}\mathscr{H}(P))$ over $X\in{\rm St}(k,n)$. With that $\widehat{P}$, settling whether $f(\widehat{P})\geq f(P)+\omega\eta$ or not can be a hard or even impossible task, for example, in Example 3.1 it is not clear if $f(\widehat{P})\geq f(P)+\omega\eta$ at all.
@@
As to the validity of the NPDo Ansatz on the objective functions in Table 1, it holds for all, except for those that involve quotients, under reasonable conditions on the constant matrices and function $\phi$. Table 2 details conditions under which the NPDo Ansatz holds, where the last column points to the places for justifications. We point out that we can take $\mathbb{P}={\rm St}(k,n)$, $Q=I_{k}$, and $\omega=1$ for all in Table 2 but judicious choices of $\mathbb{P}$ and $Q$ can increase the values of objective functions more than $\omega\eta$ as stipulated by the NPDo Ansatz for MBSub and SumCT [66, Theorem 5.2].
@@
Let $P_{*}\in{\rm St}(k,n)$ be a maximizer of (1.1). Suppose that the NPDo Ansatz holds and $P_{*}\in\mathbb{P}$. Then (2.3) holds for $P=P_{*}$ and $\Lambda=\Lambda_{*}:=P_{*}^{\operatorname{T}}\mathscr{H}(P_{*})\succeq 0$.
@@ 3.2 SCF Iteration and Convergence
3:     $\widehat{P}^{(i)}=U_{i}V_{i}^{\operatorname{T}}\in{\rm St}(k,n)$, an orthogonal polar factor of $\mathscr{H}(P^{(i)})$;
4:     calculate $Q_{i}\in{\rm St}(k,k)$ whose existence is stipulated by the NPDo Ansatz and let $P^{(i+1)}=\widehat{P}^{(i)}Q_{i}\in\mathbb{P}$;
@@
Next, we will state our convergence theorems for Algorithm 1 under the NPDo Ansatz. It is shown that as the SCF iteration progresses, the value of the objective function monotonically increases, any accumulation point of the generated approximation sequence satisfies the necessary conditions in Theorem 3.1 for a global maximizer, and under certain conditions, the accumulation point can be proved to be the limit point of the entire approximation sequence. In short, the NPDo approach is guaranteed to work.
@@
Both Theorem 3.2(b,c) have useful consequences. As a corollary of Theorem 3.2(b), we find that the NPDo Ansatz is a sufficient condition for NPDo (2.3) to have a solution because there always exists an accumulation point $P_{*}$ of the sequence $\{P^{(i)}\}_{i=0}^{\infty}$ in ${\rm St}(k,n)$.
@@
Under the NPDo Ansatz, NPDo (2.3) is solvable, i.e., there exists $P\in{\rm St}(k,n)$ such that $\Lambda=P^{\operatorname{T}}\mathscr{H}(P)\succeq 0$ and (2.3) holds.
As a corollary of Theorem 3.2(c), if $\sigma_{\min}(\mathscr{H}(P^{(i)}))$ is eventually bounded below away from 00 uniformly [^2] By which we mean that there exist a constant $\tau>0$ and an integer $K$ such that $\sigma_{\min}(\mathscr{H}(P^{(i)}))\geq\tau$ for all $i\geq K$., then
@@
    $\operatorname{rank}(\mathscr{H}(P_{*}Q))=k\quad\mbox{for any $Q\in{\rm St}(k,k)$},$ (3.7)
    then the entire sequence $\{{\cal R}(P^{(i)})\}_{i=0}^{\infty}$ converges to ${\cal R}(P_{*})$;
@@
    and if $f(P_{*})>f(P)$ for any $P\neq P_{*}$ and ${\cal R}(P)={\cal R}(P_{*})$, i.e., $f(P)$ has unique maximizer in the orbit $\{P_{*}Q\,:\,Q\in{\rm St}(k,k)\}$, then the entire sequence $\{P^{(i)}\}_{i=0}^{\infty}$ converges to $P_{*}$.
@@
In the case when objective function $f$ is right-unitarily invariant, Theorem 3.3(c) is clearly not applicable. In such a case, computing ${\cal R}(P_{*})$ is the ultimate goal because each maximizer $P_{*}$ is really a representative from the orbit $\{P_{*}Q\,:\,Q\in{\rm St}(k,k)\}$. Given $Q\in\mathbb{R}^{k\times k}$, let $g(P)=f(PQ)$. It can be verified that
@@
Thus if $f$ is right-unitarily invariant, then $g(P)\equiv f(P)$ and thus $\mathscr{H}(P)=\mathscr{H}(PQ)\,Q^{\operatorname{T}}$; if also $Q\in{\rm St}(k,k)$, then we get $\mathscr{H}(PQ)=\mathscr{H}(P)\,Q$ and as a result, condition (3.7) is equivalently to $\operatorname{rank}(\mathscr{H}(P_{*}))=k$.
@@ 3.3 Acceleration by LOCG and Convergence
Without loss of generality, let $P^{(-1)}\in{\rm St}(k,n)$ be the approximate maximizer of (1.1) from the very previous iterative step, and $P\in{\rm St}(k,n)$ the current approximate maximizer. We are now looking for the next approximate maximizer $P^{(1)}$, along the line of LOCG, according to
$P^{(1)}=\arg\max_{Y\in{\rm St}(k,n)}f(Y),\,\,\mbox{s.t.}\,\,{\cal R}(Y)\subseteq{\cal R}([P,\mathscr{R}(P),P^{(-1)}]),$ (3.9)
@@
$\mathscr{R}(P):=\operatorname{grad}f_{|{{\rm St}(k,n)}}(P)=\mathscr{H}(P)-P\cdot\frac{1}{2}\Big{[}P^{\operatorname{T}}\mathscr{H}(P)+\mathscr{H}(P)^{\operatorname{T}}P\Big{]}$ (3.10)
@@
We still have to numerically solve (3.9). For that purpose, let $W\in{\rm St}(m,n)$ be an orthonormal basis matrix of subspace ${\cal R}([P,\mathscr{R}(P),P^{(-1)}])$. Generically, $m=3k$ but $m<3k$ can happen. It can be implemented by the Gram-Schmidt orthogonalization process, starting with orthogonalizing the columns of $\mathscr{R}(P)$ against $P$ since $P\in{\rm St}(k,n)$ already. In MATLAB, to fully take advantage of its optimized functions, we simply set $W=[\mathscr{R}(P),P^{(-1)}]$ (or $W=\mathscr{R}(P)$ for the first iteration) and then do
@@
$Y=WZ\quad\mbox{for $Z\in{\rm St}(k,m)$}.$ (3.11a)
@@
$Z_{\operatorname*{opt}}=\arg\max_{Z\in{\rm St}(k,m)}\widetilde{f}(Z)\quad\mbox{with}\quad\widetilde{f}(Z):=f(WZ),$ (3.11b)
@@
~$\widetilde{\mathscr{H}}(Z)=Z\widetilde{\Lambda}\quad\mbox{with}\quad\widetilde{\Lambda}^{\operatorname{T}}=\widetilde{\Lambda}\in\mathbb{R}^{k\times k},\quad Z\in{\rm St}(k,m).$ (3.12b)
@@
Suppose that the NPDo Ansatz holds for $f$, and let $\mathbb{Z}:=W^{\operatorname{T}}\mathbb{P}\subseteq{\rm St}(k,m)$. If $W\mathbb{Z}\subseteq\mathbb{P}$, then the NPDo Ansatz holds for $\widetilde{f}$ defined in (3.11b).
@@
Let $Z\in\mathbb{Z}$ and $\widehat{Z}\in{\rm St}(k,m)$ such that
@@
Set $P=WZ\in\mathbb{P}$ (because of $W\mathbb{Z}\subseteq\mathbb{P}$) and $\widehat{P}=W\widehat{Z}\in{\rm St}(k,n)$. Noticing that ~$\widetilde{\mathscr{H}}(Z))=W^{\operatorname{T}}\mathscr{H}(WZ)$, we have (3.1) from (3.13). By the NPDo Ansatz for $f$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q=W(\widehat{Z}Q)=:W\widetilde{Z}\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$. Hence,
@@
3:     compute $W\in{\rm St}(m,n)$ such that ${\cal R}(W)={\cal R}([P^{(i)},\mathscr{R}(P^{(i)}),P^{(i-1)}])$ and $P^{(i)}$ occupies the first $k$ columns of $W$;
@@
We will perform a convergence analysis for Algorithm 2, considering an ideal situation that at its Line 4, $Z_{\operatorname*{opt}}$ is computed to be an exact maximizer of (3.11) for simplicity. We point out that the seemingly ideal situation is not completely unrealistic. In actual computation, as we explained in Remark 3.2(ii), the computed $Z_{\operatorname*{opt}}$ should be sufficiently more accurate as an approximate solution for (3.11) than $P^{(i)}$ as an approximate solution for the original problem (1.1) at that moment.